Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probably question - please explain!!

Expert replies
by gmatpup » Tue Nov 08, 2011 6:25 pm
A couple decides to have 4 children. If they succeed in having 4 children and each child is equally likely to be a boy or a girl, what is the probability that they will have exactly 2 girls and boys?

A. 3/8

B. 1/4

C. 3/16

D. 1/8

E. 1/16


Can someone please give me a step-by-step explanation as probability is my weakest area? (any probability tips would be very appreciated!)

Thanks so much!! :)
Join the discussion
Source: — Problem Solving |

by gmattaker2408 » Tue Nov 08, 2011 6:42 pm
I'm not sure. If i have to guess. I think 1/4 considering .5 probably of each girl or boy. 1/2*1/2*1/2*1/2=4/16 or 1/4

what is the correct answer?
Thanks!
Join the discussion

by gmatpup » Tue Nov 08, 2011 6:44 pm
Oops, forgot to post the answer, sorry!


Answer is: A
Join the discussion

by user123321 » Tue Nov 08, 2011 7:10 pm
if they have 4 children, the total possibilities are
2*2*2*2 = 16 possibilities
since we need to have two children out of four to be boy, we have 4c2 = 6 possibilities
so probability = 6/16 = 3/8

user123321
Just started my preparation :D
Want to do it right the first time.
Join the discussion

by gmatpup » Wed Nov 09, 2011 10:47 am
Could you pleasee elaborate on this part:

since we need to have two children out of four to be boy, we have 4c2 = 6 possibilities
so probability = 6/16 = 3/8


How did you come up with the 6 possibilities? Thanks again!!
Join the discussion

by user123321 » Wed Nov 09, 2011 11:26 am
gmatpup wrote:Could you pleasee elaborate on this part:

since we need to have two children out of four to be boy, we have 4c2 = 6 possibilities
so probability = 6/16 = 3/8


How did you come up with the 6 possibilities? Thanks again!!
out of four children, exactly two should be boys. Here order is not important. If there are two boys among the four children, then that should be fine. so it is like selecting two places among four available spaces and placing boys there.hence 4C2 = 6
(we are not required to worry about placing girls because automatically girls will come in remaining two places in 2c2 = 1 way).

user123321
Just started my preparation :D
Want to do it right the first time.
Join the discussion

by gmatpup » Wed Nov 09, 2011 11:35 am
I know this sounds silly so I apologize for asking another question but I did the 4C2 (or at least I thought I did) and I did not get 6...

Would you mind doing a step by step for me? :(
Join the discussion

by user123321 » Wed Nov 09, 2011 1:04 pm
Sure no problem.

lets use slot method. We have four blanks. each blank has two ways of filling either M(ale) or F(emale).
M/F M/F M/F M/F
so total possibilities are 2*2*2*2 = 16 possibilities.

The problem says out of those four slots, two slots should be M and other two slots should be F.
choose two slots out of four available slots and place 2M there = 4C2 = 4!/(2!.2!) = 24/4 = 6
assign remaining two slots with 2F = 2C2 = 1
so total possibilities are 6*1 = 6
so probability is 6/16 = 3/8

For your understanding I will give what those 6 possibilities are.
MMFF
MFFM
FFMM
MFMF
FMMF
FMFM

this is nothing but arranging four objects of which two are alike = 4!/2!.2! = 6

user123321
HTH.

user123321
Just started my preparation :D
Want to do it right the first time.
Join the discussion